OpenStudy (anonymous):

Fan...Medal...Testimonial...Help...Fast...Help... Solve the following inequality using an algebraic approach: -4 < 3x+5 less than or equal to 11 Can you explain how you got this answer?

OpenStudy (studygurl14):

\[-4 < 3x + 5 \le 11\] split into two \[-4 < 3x + 5\] and \[3x + 5 \le 11\] Solve, then put back together again

OpenStudy (solomonzelman):

-4 < 3x+5<= 11 -5 -5 -5 -9 < 3x <= 4 /3 /3 /3 -3 < x <= 4/3

OpenStudy (solomonzelman):

I don't really know what you mean by algebraic approach precisely.